Circular No. 6399

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744−28

R. Corbet, Goddard Space Flight Center (GSFC) and Universities Space Research Association; and K. Jahoda, GSFC, report on behalf of the RXTE Science Operations Facility and PCA team: “We continue to detect significant emission from GRO J1744−28, and the following approximate quiescent flux levels were measured in recent observations: Apr. 26, 100 mCrab, 116 min observation time; 28, 95 mCrab, 52 min; 30, 90, 100; May 3, 70, 105; 10, 60, 33; 12, 50, 58; 13, 50, 75. During these observations the following bursts were detected: Apr. 26.069 UT, 890 mCrab; 28.656, 120; May 3.785, 100; 12.688, 700. The May 12 burst had high absorption and timing and spectral characteristics typical of a type I burst. Owing to the 1^{o} PCA field of view we cannot yet completely rule out the possibility that the lower-amplitude bursts might be due to another contaminating nearby source. Pulsations at 2.1 Hz were evident during all observations, although since May 12 the pulsed fraction has decreased to below 1 percent rms. Lightcurves are available under ‘TOO News’ at

http://heasarc.gsfc.nasa.gov/0/docs/xte/xte 1st.html/.”

COMET C/1996 B2 (HYAKUTAKE) Total visual magnitude estimates: Apr. 26.12 UT, 2.2 (R. Keen, Mt.

Thorodin, CO, 7 × 50 binoculars; tail 1^{o} in p.a. 25^{o}); 26.87, 2.5, (B. H. Granslo, Tryvann, Oslo, Norway, 10 × 50 binoculars; tail 0^{o}.5 in p.a. 23^{o}); 27.86, 2.0: (O. Skilbrei, Hoenefoss, Norway, 7×50 binoculars; tail 1^{o} in p.a. 20^{o}); 28.02, 2.5–3, (J. E. Bortle, Stormville, NY, 15×80 binoculars; twilight extremely bright); May 9.82, 3.0 (G. J. Garradd, Loomberah, N.S.W., 0.25m reflector; no tail visible); 10.83, 2.8: (T. Lovejoy, Jimboomba, Queensland, 10 × 50 binoculars; tail 1^{o}); 11.81, 3.3, (Lovejoy; tail > 1^{o}); 13.81, 3.5 (Garradd, 10 × 50 binoculars; tail 3^{o}).

The first postperihelic astrometric observations were made by Garradd on May 10 and 11 and confirm the need (see IAUC 6391) to allow for nongravitational effects in the orbit solution, the following improvement being from MPEC 1996-J06. Nongravitational parameters A_{1} = +2.78 ± 0.07, A_{2} = +0.05 ± 0.08. Corrections to the ephemeris on IAUC 6391 are indicated as follows: May 12, −0^{m}.02, +0′.7; 17, −0^{m}.03, +0′.8; 22, −0^{m}.06, +1′.0; 27, −0^{m}.09, +1′.2; June 1, −0^{m}.11, +1′.3.

Epoch = 1996 Apr. 27.0 TT

T = 1996 May 1.3951 TT  ω = 130.^{o}1780
e = 0.999751  Ω = 188.0453 }2000.0
q = 0.230218 AU  i = 124.9231
1996 May 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden